ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 55P42 Kriz, Igor, University of Michigan, U. S. A. Real-oriented homotopy theory M. F. Atiyah defined a Real bundle as a complex bundle on a ℤ/2-space with anti-linear conjugation. P. S. Landweber and S. Araki used this idea to define Real cobordism and Real-oriented spectra. I will present a new Adamstype spectral sequence converging to stable homotopy groups of spheres based on Real cobordism. Connections with a ℤ/2-equivariant Adams spectral sequence based on the Eilenberg-MacLane spectrum of the constant Mackey functor ℤ/2 (extending work of J. P. C. Greenlees) will also be presented. Much of this is joint work with Po Hu. Reference: P. Hu and I. Kriz: A Real-oriented analogue of the Adams-Novikov spectral sequence, to appear